OpenStudy (anonymous):

What is the solution set for the equation 4x^2-8x+3=0?

OpenStudy (anonymous):

\[4x^{2}-8x+3=0\] Method 1: Factor a trinomial (1 - 2x)(3 - 2x) = 0 When (1 - 2x) = 0 1 - 2x = 0 x = 0.5 When (3 - 2x) = 0 2x = 3 x = 1.5 Solution x = {0.5, 1.5} Method 2: Using Quadratic Formula: \[x = \frac{ -b \pm \sqrt{b^{2} - 4ac} }{ 2a }\] where a = 4, b = -8, c = 3, \[x = \frac{ 8 \pm \sqrt{8^{2} - 4.4.3} }{ 2.4 }\] x = {0.5, 1.5}